April shoots an arrow upward into the air at a speed of 64 ft per second from a platform that is 34 feet high. The height of the arrow is given by the function h(t)=-16t^2+64t+34, where t is the time in seconds. What is the maximum height of the arrow?
Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)


Hi
The height of the arrow is given by the function h(t)=-16t^2+64t+34,
where t is the time in seconds. What is the maximum height of the arrow?
-16t^2+64t+34 |parbola opening downward, completing the square to find vertex
h(t)=-16t^2+64t+34,
h(t)=-16(t-2)^2 + 64 +34
h(t)=-16(t-2)^2 + 98 |Vertex(2,98)
98ft is the maximum height of the arrow


Answer by Alan3354(69443)

April shoots an arrow upward into the air at a speed of 64 ft per second from a platform that is 34 feet high. The height of the arrow is given by the function h(t)=-16t^2+64t+34, where t is the time in seconds. What is the maximum height of the arrow?
-------------
The max is the vertex of the parabola, on the line of symmetry t = -b/2a
t = -64/-32 = 2 seconds
h(2) = -16*4 + 64*2 + 34
h(2) = 98 feet, the max ht